# The Gödel Universe as a Lie group with left-invariant Lorentz metric and the Iwasawa decomposition

### v1

@@ 1. Introduction
Professor Karl–Hermann Neeb wrote to the author that it is possible to realize the Gödel Universe otherwise and sent an electronic version of his joint with Joachim Hilgert book [3], where in section 2.7 ’’Gödel’s cosmological model and universal covering of $\operatorname{SL}(2,\mathbb{R})$’’ is suggested a left-invariant Lorentz metric on this Lie group.
@@
In a message to the author, Professor Neeb explains this by a diffeomormism of Lie groups $\operatorname{SL}(2,\mathbb{R})$ and $\operatorname{SO}(2)\times A^{+}(\mathbb{R})$ by means of the Iwasawa decomposition for $\operatorname{SL}(2,\mathbb{R})$. Let us cite now only Theorems 6.5.1 and 9.1.3 from [6].
@@
In Theorem 3 we prove some properties of special left-invariant Lorentz metrics on three-dimensional Lie groups. Also we show in Proposition 3 that the isometry between two non-isomorphic sub-Riemannian Lie group, constructed by A. Agrachev and D. Barilari, is induced by some Iwasawa decomposition of $\operatorname{SL}(2,\mathbb{R})$.
The author expresses his gratitudes to Professor Neeb for fruitful discussions.
@@ 2. The Gödel Universe as a Lie group with left-invariant Lorentz metric
Gödel’s noticed in [1] that on $(S,ds^{2})$ acts simply transitively a four-dimensional isometry Lie group. It is easy to see that such action could be written as
@@
According to (6), the Lie subgroup $G_{3}$ can be identified with the matrix Lie group
@@ 3. The Iwasawa decompositions of Lie algebras and Lie groups
As an example, the authors of [7] give the decomposition (13) for $\mathfrak{g}=\mathfrak{sl}(n,\mathbb{R}).$ In this case $\mathfrak{k}$ is the Lie subalgebra of skew-symmetric matrices, $\mathfrak{a}$ is the Lie subalgebra of diagonal matrices (with zero trace), and $\mathfrak{n}$ is the Lie subalgebra of strictly upper triangular matrices. In particular, for $\mathfrak{g}=\mathfrak{sl}(2,\mathbb{R})$ we have
@@
The statements above imply the following Corollary.
@@
4. Applications of Iwasawa decompositions to $\mathfrak{sl}(2,\mathbb{R})$ and $\operatorname{SL}(2,\mathbb{R})$
@@
There exist an isomorphism of the Lie group $G_{3}$ onto the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ and corresponding realization of $(S_{0},ds^{2}_{0})$ as Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric.
@@
Comparing (12) and (16), we see that the linear map $\varphi:\mathfrak{a}\oplus\mathfrak{n}\rightarrow\mathfrak{g}$ such that
@@
is an isomorphism of Lie algebras. Then (18) defines isomorphism of Lie groups $\psi:\operatorname{Sol}(2)\rightarrow G_{2}:$
@@
$X=f_{0},\quad Y=-f_{1}/2,\quad Z=\sqrt{2}(f_{0}-f_{2}).$ (22)
Now we shall consider the Iwasawa decompositions for Lie groups $\operatorname{SO}(2)\times\operatorname{Sol}(2)$ and $\operatorname{SL}(2,\mathbb{R})$ and corresponding mutual expressions of partial derivatives of these Lie groups by parameters $r,s,t$ and natural left-invariant vector fields on the above Lie groups which represent some interest.
Let $kan=kan(t,s,r)$ denotes the following matrix from $\operatorname{SL}(2,\mathbb{R}):$
$\left(\begin{array}[]{cc}\cos t&\sin t\\ -\sin t&\cos t\end{array}\right)\left(\begin{array}[]{cc}e^{-s/2}&0\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}1&r\\ 0&1\end{array}\right)=$
$\left(\begin{array}[]{cc}\cos te^{-s/2}&r\cos te^{-s/2}+\sin te^{s/2}\\ -\sin te^{-s/2}&-r\sin te^{-s/2}+\cos te^{s/2}\end{array}\right).$
Then we have the following partial derivatives
$(kan)^{\prime}_{t}=\left(\begin{array}[]{cc}-\sin te^{-s/2}&-r\sin te^{-s/2}+\cos te^{s/2}\\ -\cos te^{-s/2}&-r\cos te^{-s/2}-\sin te^{s/2}\end{array}\right),$ (23)
$(kan)^{\prime}_{s}=\frac{-1}{2}\left(\begin{array}[]{cc}\cos te^{-s/2}&r\cos te^{-s/2}-\sin te^{s/2}\\ -\sin te^{-s/2}&-r\sin te^{-s/2}-\cos te^{s/2}\end{array}\right),$ (24)
$(kan)^{\prime}_{r}=\left(\begin{array}[]{cc}0&\cos te^{-s/2}\\ 0&-\sin te^{-s/2}\end{array}\right).$ (25)
The left-invariant vector fields on $\operatorname{SL}(2,\mathbb{R}),$ corresponding to $f_{0},$ $-f_{1}/2,$ and $f_{2},$ are
$v_{0}=(kan)\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)=\left(\begin{array}[]{cc}-(r\cos te^{-s/2}+\sin te^{s/2})&\cos te^{-s/2}\\ -(-r\sin te^{-s/2}+\cos te^{s/2})&-\sin te^{-s/2}\end{array}\right),$ (26)
$v_{1}=\frac{1}{2}(kan)\left(\begin{array}[]{cc}-1&0\\ 0&1\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}-\cos te^{-s/2}&r\cos te^{-s/2}+\sin te^{s/2}\\ \sin te^{-s/2}&-r\sin te^{-s/2}+\cos te^{s/2}\end{array}\right),$ (27)
$v_{2}=(kan)\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)=\left(\begin{array}[]{cc}0&\cos te^{-s/2}\\ 0&-\sin te^{-s/2}\end{array}\right)=(kan)^{\prime}_{r}.$ (28)
Comparing (23), (24), (25) and (26), (27), (28), we see that
$(kan)^{\prime}s=v_{1}-rv_{2},\quad(kan)^{\prime}_{t}=e^{-s}(v_{0}-2rv_{1}+(e^{2s}+r^{2}-1)v_{2}).$ (29)
The equalities imply that
$v_{2}=(kan)^{\prime}_{r},v_{1}=(kan)^{\prime}_{s}+r(kan)^{\prime}_{r},v_{0}=e^{s}(kan)^{\prime}_{t}+2r(kan)^{\prime}_{s}+(r^{2}+1-e^{2s})(kan)^{\prime}_{r}$ (30)
Let us consider now the Lie group $K\times AN=\operatorname{SO}(2)\times\operatorname{Sol}(2)$ with the Lie algebra $\mathfrak{k}\oplus(\mathfrak{a}+\mathfrak{n})$ and the same basis $(f_{0},-f_{1}/2,f_{2}).$ Then
$k(t)=\left(\begin{array}[]{cc}\cos t&\sin t\\ -\sin t&\cos t\end{array}\right),\quad an(s,r)=\left(\begin{array}[]{cc}e^{-s/2}&0\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}1&r\\ 0&1\end{array}\right)=\left(\begin{array}[]{cc}e^{-s/2}&re^{-s/2}\\ 0&e^{s/2}\end{array}\right),$
$k^{\prime}(t)=\left(\begin{array}[]{cc}-\sin t&\cos t\\ -\cos t&-\sin t\end{array}\right),(an)^{\prime}_{s}=\frac{-1}{2}\left(\begin{array}[]{cc}e^{-s/2}&re^{-s/2}\\ 0&-e^{s/2}\end{array}\right),(an)^{\prime}_{r}=\left(\begin{array}[]{cc}0&e^{-s/2}\\ 0&0\end{array}\right),$
$v_{0}=\left(\begin{array}[]{cc}\cos t&\sin t\\ -\sin t&\cos t\end{array}\right)\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)=\left(\begin{array}[]{cc}-\sin t&\cos t\\ -\cos t&-\sin t\end{array}\right)=k^{\prime}(t),$
$v_{1}=\frac{1}{2}\left(\begin{array}[]{cc}e^{-s/2}&re^{-s/2}\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}-1&0\\ 0&1\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}-e^{-s/2}&re^{-s/2}\\ 0&e^{s/2}\end{array}\right),$
$v_{2}=\left(\begin{array}[]{cc}e^{-s/2}&re^{-s/2}\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)=\left(\begin{array}[]{cc}0&e^{-s/2}\\ 0&0\end{array}\right)=(an)^{\prime}_{r},\quad(an)^{\prime}s=v_{1}-rv_{2},$
$k^{\prime}(t)=v_{0},\quad(an)^{\prime}s=v_{1}-rv_{2},\quad(an)^{\prime}_{r}=v_{2},$ (31)
$v_{0}=k^{\prime}(t),\quad v_{1}=(an)^{\prime}s+r(an)^{\prime}_{r},\quad v_{2}=(an)^{\prime}_{r}.$ (32)
@@
the Lorentz metric on $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ from [3] is not isometric to the subspace $(S_{0},ds_{0}^{2})$ of the Gödel Universe and the Iwasawa diffeomorphism of $\operatorname{SL}(2,\mathbb{R})$ onto $\operatorname{SO}(2)\times\operatorname{Sol}(2)$ is not isometry for Lorentz metric on $\operatorname{SL}(2,\mathbb{R})$ from [3].
@@
For any (pseudo)-Riemannian manifold $M$ with (pseudo)-metric tensor (,),($\cdot$,$\cdot$),( ⋅ , ⋅ ) , the Levi-Civita connection $\nabla$, and smooth vector fields $X,$ $Y,$ $Z$ is valid the following equation (3.5.(7)) from [8]:
@@
As a consequence, if $(M,(\cdot,\cdot))$ is a Lie group $G$ with left-invariant (pseudo)-metric (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and $X,$ $Y,$ $Z$ are left-invariant, then
@@
Let assume now that $G=\operatorname{SO}(2)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric and $(X,Y,Z)$ be an orthonormal basis of left-invariant vector fields such that $X$ be tangent to $\operatorname{SO}(2)$ and $(X,X)=1.$ Then $Y=\alpha X+Y_{1},$ $Z=\beta X+Z_{1},$ where $Y_{1},Z_{1}\in\mathfrak{sol}(2)$ and
$(\nabla_{X}Y,Z)=\frac{-1}{2}(X,[Y,Z])=\frac{-1}{2}(X,W_{1}):=-\gamma_{1},\quad W_{1}\in\mathfrak{n},$ (35)
$(\nabla_{X}Y,X)=0,\quad(\nabla_{X}Y,Y)=0,\quad\nabla_{X}Y=\gamma_{1}Z.$
We want to calculate
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y.$
@@
$(\nabla_{Y}Y,X)=(\nabla_{Y}Y,Y)=0,\quad(\nabla_{Y}Y,Z)=-(Y,W_{1}):=-\gamma_{2},\quad\nabla_{Y}Y=\gamma_{2}Z,$
$(\nabla_{X}Z,X)=(\nabla_{X}Z,Z)=0,\quad(\nabla_{X}Z,Y)=\frac{-1}{2}(X,[Z,Y])=\gamma_{1},\quad\nabla_{X}Z=-\gamma_{1}Y,$
$\quad(\nabla_{Y}Z,X)=\gamma_{1},\quad(\nabla_{Y}Z,Y)=\gamma_{2},\quad(\nabla_{Y}Z,Z)=0,\quad\nabla_{Y}Z=\gamma_{1}X-\gamma_{2}Y,$
$R(X,Y)Y=\nabla_{X}(\gamma_{2}Z)-\nabla_{Y}(\gamma_{1}Z)=-\gamma_{1}\gamma_{2}Y-\gamma_{1}^{2}X+\gamma_{1}\gamma_{2}Y=-\gamma_{1}^{2}X.$
Let us calculate yet an analogue of the sectional curvature $(R(X,Y)Y,X)=-\gamma_{1}^{2}.$
In [3], the authors consider the Lie group $\operatorname{SL}(2,\mathbb{R})$ with left-invariant Lorentz metric and orthonormal basis of opposite to above signature of the form
Then $[Y,Z]=2\sqrt{2}X$ in $\mathfrak{sl}(2,\mathbb{R}),$ but $[Y,Z]=0$ in the Lie algebra $\mathfrak{so}(2)\oplus\mathfrak{sol}(2),$ and according to the previous calcultions, $\gamma_{1}=0$ in $\operatorname{SO}(2)\times\operatorname{Sol}(2)$ and
$R(X,Y)Y=0.$ (37)
Now let us calculate $R(X,Y)Y$ and $(R(X,Y)Y,X)$ for left-invariant Lorentz metric on the Lie group $\operatorname{SO}(2)\times\operatorname{Sol}(2),$ locally isometric to the the three-dimensional subspace $(S_{0},ds_{0}^{2})$ of the Gödel Universe. According to (43) and (10), we have
$W_{1}=[Y,Z]=\sqrt{2}[-f_{1}/2,(f_{0}-f_{2})]=-\sqrt{2}f_{2},\quad\gamma_{1}=-\frac{1}{2}(X,W_{1})=\sqrt{2}.$
So, according to previous calculations,
$R(X,Y)Y=-2X,\quad(R(X,Y)Y,X)=-2.$ (38)
Let us compute now $R(X,Y)Y$ for left-invariant Lorentz metric on $\operatorname{SL}(2,\mathbb{R}),$ defined in [3]. It follows from (36) that
$[X,Y]=\sqrt{2}Z,\quad[X,Z]=\sqrt{2}Y,\quad[Y,Z]=2\sqrt{2}X.$ (39)
@@
$(\nabla_{Z}Y,X)=-2\sqrt{2},\quad(\nabla_{Z}Y,Y)=(\nabla_{Z}Y,Z)=0,\quad\nabla_{Z}Y=-2\sqrt{2}X,$
$\nabla_{Y}Z=\nabla_{Z}Y+[Y,Z]=0,$
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=-\sqrt{2}(\nabla_{Y}Z-\nabla_{Z}Y)=-4X.$ (40)
All equalities (37), (38), and (40) are pairwise different.
This implies all statements of Theorem 3. ∎

### v2

# УДК 514.82 + 512.81 + 517.977+517.838+519.46

@@ 1. Introduction
The above left-invariant Lorentz metric on $G$ and behaviour of its geodesics are defined essentially by the corresponding induced left-invariant Lorentz metric $ds^{2}_{0}$ on the subgroup $G_{3}=(\mathbb{R},+)\times\operatorname{A}^{+}(\mathbb{R}).$
Professor Karl–Hermann Neeb wrote to the author that it is possible to realize the Gödel Universe otherwise. He sent an electronic version of his joint with Joachim Hilgert book [3], where in section 2.7 ’’Gödel’s cosmological model and universal covering of $\operatorname{SL}(2,\mathbb{R})$’’ is suggested a left-invariant Lorentz metric on $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ which is isometric to $(G_{3},ds^{2}_{0})$ as stated there.
@@
In a message to the author, Professor Neeb explains the mentioned two isometries by a diffeomormism of Lie groups $\operatorname{SL}(2,\mathbb{R})$ and $\operatorname{SO}(2)\times A^{+}(\mathbb{R})$ by means of the Iwasawa decomposition for $\operatorname{SL}(2,\mathbb{R})$; let us cite Theorems 6.5.1 and 9.1.3 from [6].
@@
In Theorem 3 we prove some properties of special left-invariant Lorentz metrics on Lie groups which support the mentioned statement on the isometry of two left-invariant Lorentz metrics from [3]. Also we show in Proposition 3 that the isometry between two non-isomorphic sub-Riemannian Lie group, constructed by A. Agrachev and D. Barilari, is induced by some Iwasawa decomposition of $\operatorname{SL}(2,\mathbb{R})$.
The author expresses his gratitudes to Professor Neeb for fruitful discussions and anonymous referee for useful remarks.
@@ 2. The Gödel Universe as a Lie group with left-invariant Lorentz metric
Gödel noticed in [1] that on $(S,ds^{2})$ acts simply transitively a four-dimensional isometry Lie group. It is easy to see that such action could be written as
@@
According to (6), the Lie subgroup $G_{3}:=(\mathbb{R},+)\times G_{2}$ can be identified with the matrix Lie group
@@ 3. The Iwasawa decompositions of Lie algebras and Lie groups
As an example, the authors of [7] give the decomposition (13) for $\mathfrak{g}=\mathfrak{sl}(n,\mathbb{R}).$ In this case $\mathfrak{k}$ is the Lie subalgebra of skew-symmetric matrices, $\mathfrak{a}$ is the Lie subalgebra of diagonal matrices with zero trace, and $\mathfrak{n}$ is the Lie subalgebra of strictly upper triangular matrices. In particular, for $\mathfrak{g}=\mathfrak{sl}(2,\mathbb{R})$ we have
@@
The statements above imply the following
@@
4. $(S_{0},ds^{2}_{0})$ as (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric
@@
There exist an isomorphism of the Lie group $G_{3}$ onto the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ and corresponding realization of $(S_{0},ds^{2}_{0})$ as the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric.
@@
Comparing (12) and (16), we see that the linear map $\varphi:\mathfrak{a}\oplus\mathfrak{n}\rightarrow\mathfrak{g}_{2}$ such that
@@
is an isomorphism of Lie algebras. Let $\operatorname{Sol}(2)$ be the Lie group with the Lie algebra $\mathfrak{a}\oplus\mathfrak{n}$ for (14). Then (18) defines isomorphism of Lie groups $\psi:\operatorname{Sol}(2)\rightarrow G_{2}:$
@@
$X=f_{0},\quad Y^{\prime}=-f_{1}/2,\quad Z^{\prime}=\sqrt{2}(f_{0}-f_{2}),$ (22)
which we can change by
$X=f_{0},\quad Y=f_{1}/2,\quad Z=\sqrt{2}(f_{2}-f_{0}).$ (23)
Then
$[Y,Z]=[Y^{\prime},Z^{\prime}]=\sqrt{2}f_{2}=Z+\sqrt{2}X.$ (24)
Remark 1.
The basis (23) has a form
$X=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),Y=\frac{1}{2}\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),Z=\sqrt{2}\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right).$ (25)
The group $\operatorname{Sol}(2)$ is isomorphic to the Lie group of real lower triangular $(2\times 2)-$matrices with unit determinant.
@@
In [3], the authors consider the Lie group $\operatorname{SL}(2,\mathbb{R})$ with left-invariant Lorentz metric and orthonormal basis with the signature (,+,+)(-,+,+)( - , + , + ) of the form
We shall consider this basis as orthonormal with the signature (+,,).(+,-,-).( + , - , - ) .
@@
1) For the Lorentz metric on $\operatorname{SL}(2,\mathbb{R})$ from [3] with orthonormal basis (26) and for corresponding by the Iwasawa diffeomorphism basis on $\operatorname{SO}(2)\times\operatorname{Sol}(2)$, the curvatures $k(X,Y)=k(X,Z)=k(Y,Z)=-2,$ while for the orthonormal basis (23) on (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$, $k(X,Y)=k(X,Z)=k(Y,Z)=-\frac{1}{2}$.
2) One-parameter subgroups in $\operatorname{SL}(2,\mathbb{R}),$ defined by $X,$ $Y,$ $Z,$ are geodesics.
@@
For any (pseudo-)Riemannian manifold $M$ with (pseudo-)metric tensor (,),($\cdot$,$\cdot$),( ⋅ , ⋅ ) , the Levi-Civita connection $\nabla$, and smooth vector fields $X,$ $Y,$ $Z$ is valid the following equation (3.5.(7)) from [8]:
@@
As a consequence, if $(M,(\cdot,\cdot))$ is a Lie group $G$ with left-invariant (pseudo-)metric (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) of the signature (+,,)(+,-,-)( + , - , - ) and $X,$ $Y,$ $Z$ are left-invariant, then
@@
$[X,Y]=-\sqrt{2}Z,\quad[X,Z]=\sqrt{2}Y,\quad[Y,Z]=2\sqrt{2}X.$ (29)
@@
$(\nabla_{X}Y,Z)=0,\quad(\nabla_{X}Y,X)=(\nabla_{X}Y,Y)=0,\quad\nabla_{X}Y=0,$
$(\nabla_{Z}Y,X)=-\sqrt{2},\quad(\nabla_{Z}Y,Y)=(\nabla_{Z}Y,Z)=0,\quad\nabla_{Z}Y=-\sqrt{2}X,$
$\nabla_{Y}Z=\nabla_{Z}Y+[Y,Z]=\sqrt{2}X,$
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=\sqrt{2}\nabla_{Z}Y=-2X.$ (30)
Then $k(X,Y)=(R(X,Y)Y,X)=-2.$ Analogously, we obtain
$\nabla_{X}Z=0,\quad\nabla_{X}X=0,\quad\nabla_{Z}Z=0,\quad k(X,Z)=(R(X,Z)Z,X)=-2,$
$R(Y,Z)Z=\nabla_{Y}\nabla_{Z}Z-\nabla_{Z}\nabla_{Y}Z-\nabla_{[Y,Z]}Z=-\sqrt{2}\nabla_{Z}X=-\sqrt{2}(\nabla_{X}Z+[Z,X])=2Y,$
$k(Y,Z)=(R(Y,Z)Z,Y)=-2.$
2) $[Y,Z]=2Z+2\sqrt{2}X$ in $\mathfrak{so}(2)\oplus\mathfrak{sol}(2).$
We look for $Z$ as $\alpha f_{2}+\beta X$, see (15). It is easy to see that $\alpha=2,$ $\beta=-\sqrt{2},$
$[Y,Z]=[Y,2f_{2}-\sqrt{2}X]=2[f_{1},f_{2}]=4f_{2}=2(Z+\sqrt{2}X)=2Z+2\sqrt{2}X.$
@@
$(\nabla_{Z}Y,X)=-\sqrt{2},\quad(\nabla_{Z}Y,Y)=0,\quad(\nabla_{Z}Y,Z)=2,\quad\nabla_{Z}Y=-\sqrt{2}X-2Z,$
$\nabla_{Y}Z=\nabla_{Z}Y+[Y,Z]=\sqrt{2}X,$
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=-\sqrt{2}\nabla_{Y}Z=-2X.$
Then $k(X,Y)=(R(X,Y)Y,X)=-2.$ Analogously, we obtain
$\nabla_{X}Z=-\sqrt{2}Y,\nabla_{Z}X=\nabla_{X}Z=-\sqrt{2}Y,\quad\nabla_{Z}Z=2Y,\quad k(X,Z)=(R(X,Z)Z,X)=-2,$
$R(Y,Z)Z=\nabla_{Y}\nabla_{Z}Z-\nabla_{Z}\nabla_{Y}Z-\nabla_{[Y,Z]}Z=$
$\nabla_{Y}(2Y)-\nabla_{Z}(\sqrt{2}X)-\nabla_{(2Z+2\sqrt{2}X)}Z=$
$2Y-2\nabla_{Z}Z-2\sqrt{2}\nabla_{X}Z=2Y-4Y+4Y=2Y,$
$k(Y,Z)=(R(Y,Z)Z,Y)=-2.$
Let us compute analogous expressions for $(S_{0},ds_{0}^{2}),$ using results of Section 4.
$(\nabla_{X}Y,Z)=-\frac{\sqrt{2}}{2},\quad(\nabla_{X}Y,X)=(\nabla_{X}Y,Y)=0,\quad\nabla_{X}Y=\frac{\sqrt{2}}{2}Z,$
$(\nabla_{Z}Y,X)=-\frac{\sqrt{2}}{2},\quad(\nabla_{Z}Y,Y)=0,\quad(\nabla_{Z}Y,Z)=1,\quad\nabla_{Z}Y=-\frac{\sqrt{2}}{2}X-Z,$
$\nabla_{Y}Z=\nabla_{Z}Y+[Y,Z]=\frac{\sqrt{2}}{2}X,$
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=-\frac{\sqrt{2}}{2}\nabla_{Y}Z=-\frac{1}{2}X.$
Then $k(X,Y)=(R(X,Y)Y,X)=-\frac{1}{2}.$ Analogously, we obtain
$\nabla_{X}Z=-\frac{\sqrt{2}}{2}Y,\nabla_{Z}X=\nabla_{X}Z=-\frac{\sqrt{2}}{2}Y,\quad\nabla_{Z}Z=Y,\quad k(X,Z)=(R(X,Z)Z,X)=-\frac{1}{2},$
$R(Y,Z)Z=\nabla_{Y}\nabla_{Z}Z-\nabla_{Z}\nabla_{Y}Z-\nabla_{[Y,Z]}Z=$
$\nabla_{Y}(Y)-\nabla_{Z}\left(\frac{\sqrt{2}}{2}X\right)-\nabla_{(Z+\sqrt{2}X)}Z=$
$\frac{1}{2}Y-\nabla_{Z}Z-\sqrt{2}\nabla_{X}Z=\frac{1}{2}Y-Y+Y=\frac{1}{2}Y,$
$k(Y,Z)=(R(Y,Z)Z,Y)=-\frac{1}{2}.$
Obtained equalities imply all statements of Theorem 3. ∎
Remark 2.
It is similar, that the left-invariant Lorentz metric on $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ from [3] is isometric to Gödel metric $(S_{0},ds^{2}_{0})$ induced by Gödel metric (1) with $a=\frac{1}{\sqrt{2}}.$